7) Grandma Gertrudes Chocolates, a family owned business, has an opportunity to supply its product for distribution through a large coffee house chain. However, the coffee house chain has certain specifications regarding cacao content as it wishes to advertise the health benefits (antioxidants) of the chocolate products it sells. In order to determine the mean % cacao in its dark chocolate products. quality inspectors sample 30 pieces. They find a sample mean of 55% with a standard deviation of 4%. What is the correct t* (t critical value) used to construct a 95% confidence interval? A) 2.045 B 2.042 .699 D) 2.462 E) 2.037 C) 1.699
Answer #1

n= 30

Degree of freedom, df= n - 1 =30 - 1 =29

Significance level, \small \alpha = 1- 0.95 = 0.05

At df 29 and \small \alpha 0.05, tc for 95% confridence interval is

tc = 2.045

so answer is A) 2.045